# EventHallusion: Diagnosing Event Hallucinations in Video LLMs

### v1

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) have made significant progress in the video comprehension field. Despite remarkable content reasoning and instruction following capabilities they demonstrated, the hallucination problem of these VideoLLMs is less explored compared with its counterpart in the image domain. To mitigate this gap, we first propose EventHallusion, a novel benchmark that focuses on assessing the VideoLMMs’ hallucination phenomenon on video event comprehension. Based on the observation that existing VideoLLMs are entangled with the priors stemming from their foundation models, our EventHallusion is curated by meticulously collecting videos and annotating questions to intentionally mislead the VideoLLMs into interpreting events based on these priors rather than accurately understanding the video content. On the other hand, we also propose a simple yet effective method, called Temporal Contrastive Decoding (TCD), to tackle the hallucination problems of VideoLLMs. The proposed TCD suppresses the model’s preference toward their priors by comparing the original video with a constructed counterpart, whose temporal cues are disrupted, during the autoregressive decoding stage. Through comprehensive evaluation of eight open-source and two closed-source VideoLLMs on the proposed EventHallusion benchmark, we find that the open-source models suffer significantly from hallucination problems, whereas the closed-source models perform markedly better. By further equipping open-sourced VideoLLMs with the proposed TCD approach, evident performance improvements are achieved across most metrics in the EventHallusion benchmark. Our codes and benchmark data are available at https://github.com/Stevetich/EventHallusion.
Introduction
@@
The field of deep learning has witnessed the extraordinary advancement of Large Language Models (LLMs) (Brown et al. 2020; Touvron et al. 2023; Dettmers et al. 2024) and Multimodal Large Language Models (MLLMs) (Alayrac et al. 2022; Liu et al. 2024a; Bai et al. 2023; Jiao et al. 2024b, a). Benefiting from dramatic semantic reasoning and instruction-following capabilities, LLMs and MLLMs have become versatile assistants across multiple disciplines  (Jiao et al. 2024a; Yin et al. 2023b). Despite their great success, they still face hallucination problems, where models generate plausible yet factually incorrect responses (Leng et al. 2024), potentially threatening reliability and security in real-world applications. To systematically evaluate hallucinations of MLLMs, a thread of benchmarks has been developed.  (Li et al. 2023b; Fu et al. 2023; Liu et al. 2023b).
However, these benchmarks primarily assess hallucinations within static scenes, such as those involving objects, attributes, and inter-object relations, and therefore do not effectively address the needs of emerging VideoLLMs  (Ma et al. 2023; Li, Wang, and Jia 2023). To mitigate this gap, two recent novel benchmarks (Wang et al. 2024; Guan et al. 2024) have been proposed. Specifically, as illustrated in Fig.1, HallusionBench (Guan et al. 2024) reverses all frames within a video and then queries the model to determine whether the event described in the original captions occurs in the altered video. VideoHallucer (Wang et al. 2024) employs long videos containing multiple events and then queries the model to assess the chronological order of these events. Generally, compared to those focused on hallucinations in static images, benchmarks that assess event hallucinations in dynamic videos lag behind.
When delving into VideoLLMs’ hallucinations on events, we found that the major causes can be attributed to language priors of LLMs (Agrawal, Batra, and Parikh 2016; Agarwal, Shetty, and Fritz 2020) and statistical bias of the inherited ImageLLM’s training corpus (Leng et al. 2024). As shown in Fig.1, when presented with a woman and a bike, Video-LLaVA (Lin et al. 2023), LLaVA-NeXT-Video (Li et al. 2024a), and ShareGPT4Video (Chen et al. 2024) associate them with the intuitive assumption of ”riding the bicycle”, rather than recognizing the clear fact of ”carrying the bicycle”. However, this essential event-related hallucination cannot be effectively assessed in both VideoHallucer (Wang et al. 2024) and HallusionBench (Guan et al. 2024). As shown in Fig.1, VideoHallucer has already provided the correct events in the question, merely requiring the model to determine their chronological orders without analyzing the events themselves. Although HallusionBench addresses event-oriented hallucinations, the frame reversal operations employed to create test samples may deviate from the natural distribution of videos.
Figure 1: Comparison of our proposed EventHallusion with previous benchmarks. Different from requiring models to locate events or comprehend reversed videos, our EventHallusion focuses on evaluating inherent priors-oriented hallucinations concerning events in VideoLLMs. Green and red represent the original and hallucination settings for evaluation in HallusionBench and VideoHallucer. At the bottom, we exhibit an example in our EventHallusion where Video-LLaVA, LLaVA-NeXT-Video, and ShareGPT4Video generate hallucinations towards the video event with hallucinated content marked by red.
Toward this end, in this paper, we propose EventHallusion, a novel benchmark that systematically evaluates state-of-the-art VideoLLMs for hallucinations concerning events that are susceptible to the model’s inherent priors. The primary challenge in developing the EventHallusion is sourcing appropriate video data. Since existing publicly available video datasets, such as AcitivityNet (Caba Heilbron et al. 2015), randomly sample video fragments from daily life. As a result, their data distribution naturally reflects the priors resulting from the high-frequency co-occurrence of objects and event bindings (e.g., a person riding a bike in most cases). Therefore, instead of refactoring samples from existing video datasets, we manually collect videos and annotate question-answer pairs which may lead VideoLLMs to interpret events based on their pre-existing priors from prevalent internet video platforms. Specifically, to comprehensively probe existing VideoLLMMs’ sensitivity to their learned priors, we categorize our benchmark into two basic categories as shown in Fig.2: (1) In the first category, all videos are meticulously selected to ensure that the included events occur infrequently. Within this category, we further split videos into two groups according to their event composition. Videos in the first group (i.e., “entire rare events” in Fig.2) depict a rare event throughout the entire duration. And videos in the second group (i.e., “interleaved common-rare events” in Fig.2) include a period depicting a common event, followed by a rare event unexpectedly happens. Compared with those in the first group, videos in the second group are more challenging because the initially presented common event acts as a demonstration to consolidate the model’s tendency to their priors, which is also proved in Tab. 1. (2) In the second category, all questions are customized to include a plausible but incorrect common event as a distraction, and videos depict events that occur frequently. Instead of evaluating VideoLLM’s sensitivity to video content as in the first category, the testing samples in the second category are designed to assess the model’s sensitivity to question content, where the included misleading event descriptions can trigger the model to rely on its pre-existing priors. Through comprehensive evaluation, we find that existing open-source VideoLLMs suffer from severe event hallucination problems on our benchmark, which can provide valuable insights for further research.
Moreover, inspired by a hallucination elimination approach for image LLMs (i.e., VCD (Leng et al. 2024)), we propose a simple yet effective method called Temporal Contrastive Decoding (TCD) to reduce the event-related hallucinations of existing VideoLLMs. As previously analyzed, VideoLLMs can be easily influenced by their prior knowledge. Therefore, our TCD is designed to mitigate such biases and enhance the model’s objective judgments in event comprehension. Technically, for each given video, we first construct its counterpart with temporal downsampling and shuffling data augmentation operations. Both the original video and its counterpart are then fed into the VideoLLMs simultaneously to generate a pair of predicted logits, which are used for contrastive decoding similar to the approach in VCD (Leng et al. 2024). In the proposed TCD, the created counterpart disrupts the temporal cues while preserving spatial object and scene information, making its predicted logits more susceptible to bias from the model’s prior knowledge. Subtracting these amplified priors in the logits space during the contrastive decoding process helps guide the model toward more reliable judgments. As a plug-and-play approach, our TCD can consistently boost existing VideoLLM’s performances on the proposed EvenHallusion as shown in Tab.2.
In summary, our contributions are threefold: (1) We propose EventHallusion, a novel video QA benchmark focused on assessing event-related hallucinations in VideoLLMs; (2) We introduce Temporal Contrastive Decoding (TCD), a simple yet effective approach that can reduce hallucinations across a wide range of VideoLLMs in a plug-and-play manner; (3) We conduct comprehensive ablation studies to provide more in-depth analysis.
@@
Figure 2: Illustration of evaluation categories. Our EventHallusion is basically divided into two categories: Sensitivity to Video Content and Sensitivity to Question Content. The content that induces VideoLLMs to generate event-related hallucinations is marked by red.
@@
Related Works
Video-supported MLLMs are further proposed since the established vision-text understanding ability of image MLLMs. To enable MLLMs to comprehend video content and the concept of temporality, numerous methods have been put forward. Video-ChatGPT (Maaz et al. 2023) utilizes spatial and temporal pooling operations to conduct video modeling, while Video-LLaMA (Zhang, Li, and Bing 2023) introduces a Video Q-Former to aggregate frame-level representations. Vista-LLaMA (Ma et al. 2023) preserves the equal distance between all visual tokens and any language tokens and proposes the temporal Q-Former projector to further enhance the representations of the video and the modeling of temporality. LLaMA-VID (Li, Wang, and Jia 2023) proposes a dual-token paradigm, using a context token and a content token to represent each frame, which effectively supports the comprehension of long videos and maintains temporal information. However, despite their excellent performance on existing benchmarks, these methods still suffer the hallucination problem in temporality.
Video Evaluation Benchmarks. Despite the extraordinary advancement of MLLMs, it is of great significance to conduct quantitative evaluations of existing models. To thoroughly evaluate the performance of image MLLMs, a wide scope of efforts has been made (Fu et al. 2023; Liu et al. 2023b; Li et al. 2024c). Their evaluation comprises various aspects, including perception, reasoning, and cognition. With the development of VideoLLMs, several benchmarks have also been proposed to assess video comprehension capacities. Tempcompass (Liu et al. 2024b) focuses on the thorough evaluation of the temporal perception ability of VideoLLMs, including action, speed, direction, attribute change, and event order. FunQA (Xie et al. 2023) proposes utilizing humorous, creative, and magic videos to assess the in-depth comprehension ability of existing VideoLLMs.
Hallucination in MLLMs. Hallucination is a phenomenon widely observed in MLLMs. The occurrence of hallucinations is attributed to various factors, including imbalanced training data distribution, weak attention to visual content, and other possible reasons. To alleviate the impact of hallucinations in image MLLMs, several methods have been proposed. Woodpecker (Yin et al. 2023a) adopts a post-hoc correction mechanism that utilizes additional extracted visual cues to rectify previously generated responses, while Volcano (Lee et al. 2023) proposes a critique-revise-decide process to self-revise their initial generated answers. Besides, LRV-Instruction (Liu et al. 2023a) and HalluciDoctor (Yu et al. 2024) propose to mitigate this problem from the data perspective to provide more balanced, high-quality instruction tuning datasets.
Some other works propose to alleviate hallucination during the inference phase, in which the contrastive decoding (CD) method has been widely adopted due to its excellent performance and the absence of additional training. VCD (Leng et al. 2024) employs the original and distorted images for comparison in contrastive decoding, thereby reducing the impact of statistical bias and language priors, while IBD (Zhu et al. 2024) constructs the comparison pair by introducing the image-biased and original MLLMs. Despite the extensive exploration of adopting CD in ImageLLMs, there is still a lack of effort in employing CD to handle hallucinations in VideoLLMs. Inspired by this, we propose Temporal Contrastive Decoding (TCD) to mitigate event hallucinations in VideoLLMs.
EventHallusion Benchmark
In this section, we first analyze the vulnerabilities of VideoLLMs emphasized by each category in our EventHallusion benchmark. Then, we elaborate on the detailed curation process of video data and question-answer pair annotation. Finally, we introduce the metrics calculation approach.
Vulnerability Dimension Analysis
To comprehensively assess a VideoLLM’s sensitivity to its priors across diverse aspects, we employ two categories to examine the influence of video and question content, respectively.
Sensitivity to Video Content
In this category, we intentionally collect videos depicting events that infrequently occur in daily life (e.g., a person carrying a bike, as shown in Fig.1). Such videos can easily trigger the prior knowledge learned from foundational ImageLLMs (noting that existing open-source VideoLLMs are all derived from certain ImageLLMs). As a result, the VideoLLM may rely on these strong visual priors rather than accurately interpreting the temporal cues of the video.
Sensitivity to Question Content
In this category, we select videos depicting events that commonly occur in daily life to exclude the influence of the visual priors mentioned previously. Additionally, we introduce a frequently occurring event that is inconsistent with the ground truth depicted in the video (e.g., “eating” in the Fig.2). Such customized questions can easily trigger priors of the LLM which obtains knowledge from enormous textual corpus. As a result, the VideoLLM may rely on its language priors to find a shortcut in answering the question.
Video Data Collection
When collecting video sources for our EventHallusion, we adhere to the following two rules to ensure high quality. Firstly, the objects and inter-object interaction details should be presented clearly to avoid visual detail missing. Secondly, the scene presented in the video is kept consistent without any shot cuts to avoid the adverse effects of environmental variation.
In total, we collect 400 eligible videos from multiple open video platforms on the Internet. To avoid data leakage, we have blurred the captions in the video, thereby encouraging models to analyze video events solely through visual content. These 400 videos cover 7 domains including Pet&Animal, Sports Competition, Food&Drink, Gym Exercises, Vehicle, Life Record, and Nature as shown in Fig.4. And the detailed duration distribution of these videos is shown in Fig.5.
QA Pairs Annotation
When annotating the questions and answers for the collected videos, we devise two types of QA pair designs. Firstly, we formulate the QA pair as a binary classification task, where the question asks whether a certain event occurs in the given video, and the model is required to make a choice between “yes” and “no”. Besides, we also ask the model to freely describe the video contents in an open-ended manner. And then we match the detailed description with the ground-truth events.
Evaluation Metric Calculation
Binary Classification.
Since we have formulated the answer as either ”yes” or ”no”, we simply count the first word of the generated response and compare it with the ground-truth answer. The accuracy is provided for every category respectively.
Detailed Description Matching.
For each generated detailed description, we employ the powerful GPT-4o to examine whether hallucinations have occurred by comparing generated responses with ground-truth descriptions. More details about templates for prompting GPT-4o can be seen in the Appendix.
Figure 4: Video category and questions distribution. “Entire” implies that the entire video depicts the rare event. “Interleave” implies that common and rare events interleavedly occur in the video. “Misleading” implies the video with the common event, but with its question customized to include misleading event demonstration.
Figure 5: Video duration distribution
Temporal Contrastive Decoding
In this section, we propose Temporal Contrastive Decoding (TCD) to alleviate the event hallucination in VideoLLMs. We first explore the influence of sampled frame numbers on the accuracy and then introduce details about our proposed TCD method.
Impact of Sampled Frame Numbers
While current VideoLLMs exhibit a universal deficiency in dealing with event hallucination problems, their accuracy on our binary classification task can be influenced by the number of sampled frames. We believe that more sampled frames can enhance the temporal modeling of events, thereby improving the comprehension of VideoLLMs. To verify our hypothesis, we conduct the experiment to explore the overall accuracy of VideoLLMs on EventHallusion with different sampled frame numbers. As results have shown in Figure 6, the overall accuracy of LLaVA-Next (Li et al. 2024a) and Vila (Lin et al. 2024) increases when more sampled frames are provided within a certain range. On the contrary, fewer sampled frames can compromise temporal information and cause event hallucinations.
Figure 6: Overall accuracy on binary classification tasks with different sampled frame numbers.
Contrast with distorted temporality
The above analysis indicates that sampled frame numbers can influence the modeling of temporal information in VideoLLMs, which inspires us to adopt Contrastive Decoding (CD) (Li et al. 2022) to mitigate the event hallucination phenomenon. Given a textual input $x$ and the video $V$, the auto-regressive decoding process in VideoLLMs can be formulated as follows:
$p\left(y_{t}\mid S(V),x,y_{<t};\theta\right)=\operatorname{softmax}\left(\operatorname{logit}_{\theta}\left(y_{t}\mid S(V),x,y_{<t}\right)\right)$ (1)

### v2

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) have made significant progress in the video comprehension field. Despite remarkable content reasoning and instruction following capabilities they demonstrated, the hallucination problem of these VideoLLMs is less explored compared with its counterpart in the image domain. To mitigate this gap, we propose EventHallusion, a novel benchmark that focuses on assessing the VideoLLMs’ hallucination toward event, the crux of video analysis. From a hallucination attribution perspective, our EventHallusion benchmark is curated to assess a VideoLLM’s susceptibility toward language priors and vision-language biases. On the other hand, we also propose a simple yet effective method, called Temporal Contrastive Decoding (TCD), to tackle the hallucination problems of VideoLLMs. The proposed TCD method rectifies the model’s bias toward its priors during the decoding stage by comparing the original video with a modified version, in which temporal cues are disrupted. Through comprehensive evaluation of eight open-source and two closed-source VideoLLMs on the proposed EventHallusion benchmark, we observe that the open-source models suffer significantly from hallucination problems, whereas the closed-source ones perform markedly better. By further equipping open-source VideoLLMs with the proposed TCD approach, evident performance improvements are achieved across most metrics in the EventHallusion benchmark. Our codes and benchmark data are available at https://github.com/Stevetich/EventHallusion.
Figure 1: Illustration of two types of underlying reasons that may lead to existing VideoLLMs’ hallucinations. We use red to indicate the hallucinations in answers generated by existing state-of-the-art VideoLLMs, namely Video-LLaVA and ShareGPT4Video.
1 Introduction
@@
The field of deep learning has witnessed the extraordinary advancement of Large Language Models (LLMs) [6, 36, 10] and Multimodal Large Language Models (MLLMs) [4, 28, 5, 15, 13]. Benefiting from dramatic semantic reasoning and instruction-following capabilities, LLMs and MLLMs have become versatile assistants across multiple disciplines  [13, 41]. Despite their great success, they are still trapped in the hallucination problem, where models generate plausible but factually incorrect responses [18], potentially threatening reliability and security in applying them in real-world scenarios [33, 14]. To systematically evaluate hallucinations of MLLMs, a thread of benchmarks has been developed.  [22, 11, 29]. Specifically, POPE examines dataset distributions to investigate the hallucination phenomenon in MLLMs, focusing on frequently occurring objects and common object combinations, while MME and MMBench assess hallucinations through systematic evaluations of object existence, attributes, and relations.
However, these benchmarks primarily assess hallucinations within static scenes, such as those involving objects, attributes, and inter-object relations, and therefore do not effectively address the needs of emerging VideoLLMs  [31, 23]. Two recent benchmarks [37, 12] further extend their evaluation to dynamic scenes with event-related queries. In pursuit of the rigorous evaluation, both of them curate more challenging event comprehension tasks. Specifically, HallusionBench [12] reverses all frames within a video, subsequently querying the model to determine whether the original event occurs in the altered video. VideoHallucer [37] selects long videos containing multiple events and then queries the model to evaluate the chronological order of these events. Although advancing the evaluation of event hallucination, they still suffer from the following limitations. On the one hand, the frame-reversed video in HallusionBench can deviate from the distribution of natural videos, as not all events remain meaningful when played in reverse. On the other hand, merely judging the chronological order of events given in the questions fails to capture the model’s authentic ability to comprehend the detailed content of the events [^1] A more detailed analysis of event-related VQA samples in HallusionBench [12] and VideoHallucer [37] is provided in the supplementary materials due to page limitations..
When delving into VideoLLMs’ event-related hallucinations, we identify two primary causes rooted in the development processes of prevalent VideoLLMs. Firstly, building upon the foundation of LLMs, VideoLLMs inherently inherit their language priors [3, 2], which manifests as an overconfidence in their language knowledge. As illustrated in the left column of Fig.1, both Video-LLaVA and ShareGPT4Video are easily misled by the frequently occurring but misleading event “person drink coffee” asked in the question. Moreover, due to the high cost associated with scaling up high-quality video-language data, prevalent VideoLLMs typically acquire multimodal knowledge from large-scale image-language datasets. However, in these datasets, specious event-related captions often arise based solely on a single image, which inevitably introduces spurious scene-event correlations. As shown in the right column of Fig.1, both Video-LLaVA and ShareGPT4Video intuitively make the conclusion of “cats are playing” when perceiving two cats within the video, without respecting the truth.
Toward this end, in this paper, we propose EventHallusion, a novel benchmark that systematically evaluates event-related hallucinations of state-of-the-art VideoLLMs. From a hallucination attribution perspective, our EventHallusion is curated to assess the susceptibility of a VideoLLM to language priors and vision-language correlation biases, respectively. Specifically, our benchmark consists of two basic categories as shown in Fig.2: (1) Susceptibility to language priors: In this category, all questions are annotated with a misleading event designed to prompt the model to make decisions based on its language priors. Within such events, while objects align with those depicted in the video, their actions are deliberately misrepresented. (2) Susceptibility to vision-language correlation biases: In this category, all videos are meticulously selected to ensure that the included events are infrequent yet plausible. When presented with these infrequent events, the model can easily be misled to interpret the event as one that frequently co-occurs with the objects or scenes depicted in the video, based on its memory. It is worth mentioning that through comprehensive evaluation, we find that existing open-source VideoLLMs suffer from severe event hallucination issues on our benchmark, which could offer promising avenues for enhancing the VideoLLM’s capabilities in the future research.
@@
Figure 2: Illustration of examples in our EventHallusion benchmark. Questions are basically divided into two categories: Susceptibility to Language Prior and Susceptibility to Vision-Language Correlation Bias. The contents that induces VideoLLMs to generate event-related hallucinations are marked by red.
Moreover, we propose a simple yet effective method called Temporal Contrastive Decoding (TCD) to reduce the event-related hallucinations of existing VideoLLMs. Without any additional training budgets, our TCD rectifies the model’s bias to its prior during the decoding stage. Technically, for each given video, we first apply the temporal downsampling operation to create a modified counterpart. Both the original video and its modified counterpart are then input into the VideoLLMs. During the decoding stage, the logits prediction of the original video is adjusted by the logits prediction of the counterpart. The rectified logits are then used to sample the next token and continue the decoding process. Within the proposed TCD, the modified counterpart disrupts the temporal cues while preserving spatial object and scene information, making its predicted logits more susceptible to bias from the model’s event-related prior knowledge. Subtracting these amplified priors in the logits space during the decoding process helps guide the model to make more reliable judgments. As a plug-and-play approach, our TCD can consistently boost existing VideoLLM’s performances on the proposed EvenHallusion as shown in Tab.2.
In summary, our contributions are threefold: (1) We propose EventHallusion, a novel video QA benchmark focused on assessing event-related hallucinations in VideoLLMs; (2) We introduce Temporal Contrastive Decoding (TCD), a simple yet effective approach that can reduce hallucinations across a wide range of VideoLLMs in a plug-and-play manner; (3) We also conduct comprehensive ablation studies to provide more in-depth analysis.
2 Related Works
Video-supported MLLMs are further proposed since the established vision-text understanding ability of image MLLMs. To enable MLLMs to comprehend video content and the concept of temporality, numerous methods have been put forward. Video-ChatGPT [32] utilizes spatial and temporal pooling operations to conduct video modeling, while Video-LLaMA [43] introduces a Video Q-Former to aggregate frame-level representations. Vista-LLaMA [31] preserves the equal distance between all visual tokens and any language tokens and proposes the temporal Q-Former projector to further enhance the representations of the video and the modeling of temporality. LLaMA-VID [23] proposes a dual-token paradigm, using a context token and a content token to represent each frame, which effectively supports the comprehension of long videos and maintains temporal information. However, despite their excellent performance on existing benchmarks, these methods still suffer the hallucination problem in temporality.
Video Evaluation Benchmarks. Despite the extraordinary advancement of MLLMs, it is of great significance to conduct quantitative evaluations of existing models. To thoroughly evaluate the performance of image MLLMs, a wide scope of efforts has been made [11, 29, 24]. Their evaluation comprises various aspects, including perception, reasoning, and cognition. With the development of VideoLLMs, several benchmarks have also been proposed to assess video comprehension capacities. Tempcompass [30] focuses on the thorough evaluation of the temporal perception ability of VideoLLMs, including action, speed, direction, attribute change, and event order. FunQA [38] proposes utilizing humorous, creative, and magic videos to assess the in-depth comprehension ability of existing VideoLLMs.
Hallucination in MLLMs. Hallucination is a phenomenon widely observed in MLLMs. The occurrence of hallucinations is attributed to various factors, including imbalanced training data distribution, weak attention to visual content, and other possible reasons. To alleviate the impact of hallucinations in image MLLMs, several methods have been proposed. Woodpecker [40] adopts a post-hoc correction mechanism that utilizes additional extracted visual cues to rectify previously generated responses, while Volcano [17] proposes a critique-revise-decide process to self-revise their initial generated answers. Besides, LRV-Instruction [27] and HalluciDoctor [42] propose to mitigate this problem from the data perspective to provide more balanced, high-quality instruction tuning datasets.
Some other works propose to alleviate hallucination during the inference phase, in which the contrastive decoding (CD) method has been widely adopted due to its excellent performance and the absence of additional training. VCD [18] employs the original and distorted images for comparison in contrastive decoding, thereby reducing the impact of statistical bias and language priors, while IBD [44] constructs the comparison pair by introducing the image-biased and original MLLMs. Despite the extensive exploration of adopting CD in ImageLLMs, there is still a lack of effort in employing CD to handle hallucinations in VideoLLMs. Inspired by this, we propose Temporal Contrastive Decoding (TCD) to mitigate event hallucinations in VideoLLMs.
3 EventHallusion Benchmark
In this section, we first elaborate on the data curation process of two basic categories of our EventHallusion benchmark. Afterward, we demonstrate the distribution statistics of the whole EventHallusion benchmark. Finally, we introduce the specific evaluation protocol of different question types.
3.1 Data Curation
Susceptibility to language priors. In this category, we first collect videos describing the events that occur in daily life from the internet. To avoid information leakage, we blur the text captions printed on videos, thereby encouraging models to analyze events solely based on visual content. However, when annotating the questions, we explicitly include another frequently occurring event that is inconsistent with the video content to trigger the language priors within the LLM. With these “misleading triggers”, we formulate questions as a binary choice problem as shown below:
“Did the [subject] [event] happen in the video?”
Here, the combination of [subject] and [event] correspond to the “misleading triggers” aforementioned. In Fig.2, we refer to this type of sample as “Common Events with Misleading Questions”.
Susceptibility to vision-language correlation biases. In this category, the primary challenge lies in sourcing the appropriate video data. Since existing publicly available video datasets (e.g, ActivityNet [7], [9], [34]) collect videos from daily life, the events they include cater to the vision-language correlation biases stored in VideoLLMs, resulting in high recognition precision. Therefore, instead of refactoring existing video datasets, we meticulously collect videos from prevalent Internet video platforms and filter them with rigorous cross-validation. Specifically, we begin by manually collecting videos that depict infrequent events based on human consensus. To minimize personal bias, we engage three annotators to assess whether the events in these selected videos are indeed uncommon. Only videos that receive unanimous positive judgments from all annotators are retained as testing samples. Similar to the first category, we also employ the blurring operation on videos to prevent information leakage. We further divide these videos into two groups: (1) Videos where infrequent events occur throughout the entire duration, referred to as “Entire Rare Events” in Fig.2. (2) Videos containing a mix of both frequent and infrequent events, referred to as “Mix Common-Rare Events” in Fig.2.
For constructing questions, we design both binary and open-ended types for comprehensive evaluation. Specifically, the binary questions ask the model whether a plausible but incorrect common event occurs in the given video. The open-ended questions, on the other hand, simply prompt the model to freely describe the video content. Formally, binary and open-ended questions for the “Entire” category can be formulated as presented below:
Binary: “Did the [subject] [event] in the video?”,
Open-Ended: “Please describe this video in detail.”,
For the “Mix” category, since both common and rare events can happen in the video, we reformulate the binary questions as below:
“Did the [subject] [event] throughout the entire video?”,
“Did any accident or anything unexpected happen in the video?”,
Dataset distribution statistics. In total, our EventHallusion benchmark comprises 400 videos and 711 questions. To ensure comprehensive coverage across diverse domains, these videos span seven scenarios, including Pet&Animal, Sports Competition, Food&Drink, Gym Exercises, Vehicle, Life Record, Nature, as shown in Fig.3. The distribution of video lengths in our benchmark is presented in Fig.4, showing that most videos are relatively short. This is because longer videos often involve significant scene fluctuations, which introduce additional challenges for VideoLLMs beyond the uni-modal and multi-modal biases we aim to examine in this paper.
Figure 3: Video category and questions distribution. “Entire” implies that the entire video depicts the rare event. “Mix” implies that common and rare events interleavedly occur in the video. “Misleading” implies the video with the common event, but with its question customized to include misleading event demonstration.
Figure 4: The distribution of video length within our EventHallusion benchmark.
3.2 Evaluation Protocol
Binary question answering. Following previous benchmarks [37, 22], we compare the first word of the model’s answer with the ground-truth answer to determine whether the model answered the question correctly. This approach is justified by the VideoLLM’s exceptional instruction-following capability, which typically enables it to answer binary choice questions with “yes” or “no” at the beginning. We calculate the accuracy as the evaluation metric for binary questions.
Open-ended question answering. For automatic evaluation, we feed both generated and ground-truth answers into GPT-4o, using it as a knowledgeable evaluator to determine whether the model answers the question correctly following prior benchmarks [30]. We also use accuracy as the evaluation metric. Detailed templates of prompting GPT-4o are provided in the supplementary materials due to page limitation.
@@ 3.2 Evaluation Protocol
4 Temporal Contrastive Decoding
To reduce the aforementioned hallucinations in VideoLLMs, the most straightforward approach is to further enrich debiased data for finetuning. However, this approach is particularly resource-intensive for collecting such videos and annotating QA pairs. To address this, we present our Temporal Contrastive Decoding (TCD) method, which effectively alleviates hallucinations in videos without requiring additional training efforts.
Motivated by the previous work [18], our TCD method aims to reduce event hallucinations by adjusting the distribution of predicted logits at each decoding step, using a contrastive counterpart that is more aligned with the model’s inherent biases. For clarity, we first review the standard auto-regressive decoding process of a VideoLLM. Given the text instruction $x$ and the video $V$, the predicted logit $z_{t}^{ori}$ at the $t$-th timestep can be formulated as:
$z_{t}^{ori}=\operatorname{logit}_{\theta}\left(y_{t}\mid V,x,y_{<t}\right)$ (1)